# DevBench: A multimodal developmental benchmark for language learning

### v1

@@ Abstract
How (dis)similar are the learning trajectories of vision–language models and children? Recent modeling work has attempted to understand the gap between models’ and humans’ data efficiency by constructing models trained on less data, especially multimodal naturalistic data. However, such models are often evaluated on adult-level benchmarks, with limited breadth in language abilities tested, and without direct comparison to behavioral data. We introduce DevBench, a multimodal benchmark comprising seven language evaluation tasks spanning the domains of lexical, syntactic, and semantic ability, with behavioral data from both children and adults. We evaluate a set of vision–language models on these tasks, comparing models and humans not only on accuracy but on their response patterns. Across tasks, models exhibit variation in their closeness to human response patterns, and models that perform better on a task also more closely resemble human behavioral responses. We also examine the developmental trajectory of OpenCLIP over training, finding that greater training results in closer approximations to adult response patterns. DevBench thus provides a benchmark for comparing models to human language development. These comparisons highlight ways in which model and human language learning processes diverge, providing insight into entry points for improving language models.
@@ 1 Introduction
Recent work has attempted to bridge this data gap by constructing models that are trained on less data, including on naturalistic data from children [2, 3, 4, 5]. However, these models have typically been evaluated using benchmarks that reflect adult human performance, whether explicitly (i.e., with comparison data collected from adult participants) or implicitly (i.e., high accuracy on some recognition task). Such evaluation methods are not appropriate to the goal of understanding whether developmentally realistic data leads to human-like learning. We would not expect child performance to be similar to adult performance on these tasks for various reasons either related to language competence (i.e., vocabulary knowledge) or more generally to co-developing cognitive abilities (e.g., working memory limitations and vocabulary knowledge). Instead, it is crucial to evaluate models on benchmarks that can indicate whether the language ability gained by machine learning models matches the language ability gained by children when exposed to similar developmental data.
@@
Additionally, to conduct methodologically rigorous cognitive evaluation of machine learning models, it is important to compare human and model performance directly (rather than assuming that humans will perform well), and to ensure that evaluation conditions are as similar as possible for models and humans [11, 12]. Notably, many language evaluations with infants and toddlers are multimodal in nature, because this method permits the measurement of language comprehension without the additional working memory and motor control demands of language production. Young children can use nonverbal response methods such as looking or pointing, which can nonetheless reflect their language abilities. The incorporation of multimodality has an additional advantage of introducing grounding, which has been suggested as a possible solution to the data gap [1].
@@
To create a benchmark that satisfies these desiderata, we introduce DevBench, a multimodal benchmark of language evaluation tasks. DevBench includes a suite of seven tasks measuring lexical, syntactic, and semantic ability, with human data from both children and adults. Notably, the primary evaluation metric reflects models’ similarity to human response patterns, rather than absolute performance, allowing us to capture finer-grained details about human–model similarity.
@@ 2.1 Multimodal benchmarks
Since the advent of multimodal models, various benchmarks have been developed to evaluate their performance [14]. A majority of these benchmarks primarily involve visual question answering [15, 16, 17, 18, 19]; other tasks include image captioning [16, 20], prediction [21], and retrieval [15, 20].
@@ 2.2 Developmentally inspired evaluation
Some recent work in machine learning evaluation has used benchmarks inspired by developmental psychology and cognitive science. For example, the Large Language Model Response Score (LRS) [22] draws from key experiments in the child development literature to construct question answering tasks for large language models, though it does not involve comparison to data from children. In addition, most of its tasks were originally multimodal (with visual stimuli and verbal prompts), and it is not clear how the translation into a unimodal verbal task would affect human performance. The Machine Word Learning (MEWL) benchmark [23] is a multimodal evaluation that builds upon hypothesised mechanisms of few-shot word learning in children, and requires the inference of the meaning of novel words from a set of visual scenes with referentially ambiguous labels. The benchmark contains data from adults, but their poor performance on some subsets of these trials (e.g., those requiring pragmatic implicature) suggests that children might find these tasks very difficult.
@@ 3 DevBench description
DevBench contains seven tasks across lexical, syntactic, and semantic domains. Each task is accompanied by item-level human data so that full human response distributions can be compared to model scores. The lexical tasks measure vocabulary knowledge, operationalised as the ability to correctly pick out the visual referent of a noun label. The syntactic tasks measure grammatical knowledge, operationalised as the ability to correctly pick out a scene containing the correct relations among its constituent members when given a sentential label. The semantic tasks measure conceptual (visual or linguistic) representation space via representational similarity. A schematic of the tasks and the ages of participants contributing data for each task is shown in Figure 1, and sample trials are shown in Figure 2. All code and data will be available at github.com/alvinwmtan/dev-bench once our unpublished datasets have been cleared for dissemination; all data are licensed under CC BY-NC-SA or more permissible licences, and have been anonymised to eliminate any personally identifiable information.
@@ 3.1 Tasks
In this task, children are presented with two visual images (one target and one distractor) as well as a verbal cue (e.g., “Look at the dog!”) [29]. Children’s proportion looking time to each image is measured. We used data from three datasets to improve age coverage: 18-month-old data from Adams et al. [30], 24-month-old data from Frank et al. [31], and 30-month-old data from Donnelly and Kidd [32] (total $N$ = 197), with replacement stimuli for images that were not sharable.
@@
In this task, participants were presented with a cue word, and asked to name the first word association that came to mind. We use association frequencies as a proxy for similarity scores. We combined a child dataset from Entwisle [38] with a subset of the adult dataset from Nelson et al. [39] comprising all cue words found in the child dataset (total $N$ = 1189), and thresholded the included responses to omit idiosyncratic responses.
@@
In this task, 4-, 10-, and 14-month-old infants saw pairs of images, and proportion looking times to each image was measured [40] (total $N$ = 73). We used replacement stimuli for some images that were not sharable in the original stimuli set. Dissimilarities between images were calculated as the difference in proportion looking times to the two images.
@@ 3.2 Human–model comparison
On lexicon and syntax tasks, our goal was to compare the distribution of human choices across response options to model choices. We obtained image–text matching logits for each response option, and calculated the softmax-optimised Kullback–Leibler divergence of human responses from model responses. This was operationalised as the minimum KL divergence between the human response probability distributions $h$ from model logits $m$, optimizing the softmax exponent $\beta$. We average this divergence across trials $t$, and calculate each distribution across images $i$:
@@
For the visual semantic tasks, we conducted human–model comparison by applying representational similarity analysis (RSA) [44] on human and model representational dissimilarity matrices. For WAT, we calculated the softmax-optimised KL divergence of human association probabilities from softmaxed model text embedding similarities, averaged over all trials. All comparisons were conducted within each age bin when applicable.
@@ 4 Benchmark
We evaluated a diverse set of vision–language models on our benchmark, using only CPUs to run inference on existing models. A description of all models can be found in the SI. A summary of model performance for each task (averaged across all ages) is shown in Table 2, along with model characteristics (number of parameters and size of training set). For lexicon and syntax tasks as well as WAT, we report $D^{*}_{\textup{KL}}$, for which lower scores indicate greater human-likeness in response patterns. For visual semantic tasks (VOC, THINGS), we report RSA similarity, for which higher scores indicate greater human-likeness in response patterns. Divergences are not comparable across datasets due to different features of each dataset.
Table 2: Model characteristics and performance across all tasks, demonstrating variation across models. Arrows indicate the direction of better performance (i.e., lower is better vs. higher is better). Bolded results indicate best result on a task.
@@
Overall, CLIP-large and OpenCLIP-H performed relatively well on the lexicon and syntax tasks. CVCL, which was trained on a small set of annotated head-mounted camera from infants [4], was mostly dissimilar to humans in lexicon and syntax tasks, even on the looking-while-listening (LWL) task, which was administered to young children. However, CVCL outperformed other models on the visual object categorization task, suggesting that its visual representational space was more similar to infants’.
@@ 6.1 Limitations and future work
The development of DevBench was constrained by currently available human data; notably, this means that some tasks have relatively few items (in comparison with many other machine learning benchmarks), and some tasks have relatively few human participants. These limitations may result in uncertain reliabilities for the obtained results. More data are currently being collected for some of the tasks in DevBench, which we anticipate will improve reliability, and simultaneously permit more expansive developmental analyses due to an extension of the included age ranges. However, we hope that this approach and standardized format will enable future researchers to contribute novel datasets to DevBench, which we anticipate may grow in the coming years. Indeed, DevBench also includes only tasks for English-speaking children and adults, due to dataset availability. This situation reflects inequalities in language acquisition research [52], and more data is needed to construct a multilingual version of DevBench, which will be more comprehensive and generalisable.
@@ 6.2 Conclusion
We hope that DevBench can serve as an encouragement for machine learning researchers and cognitive scientists to develop vision–language models that not only perform well, but also can more closely approximate human learning. This correspondence should be pursued both as a way to better understand the processes underlying human development and as a way to transfer humans’ learning efficiencies onto models.
@@
Appendix A Appendix
A.1 Evaluated models
@@
A.2 Detailed benchmark results

### v2

@@ Abstract
How (dis)similar are the learning trajectories of vision–language models and children? Recent modeling work has attempted to understand the gap between models’ and humans’ data efficiency by constructing models trained on less data, especially multimodal naturalistic data. However, such models are often evaluated on adult-level benchmarks, with limited breadth in language abilities tested, and without direct comparison to behavioral data. We introduce DevBench, a multimodal benchmark comprising seven language evaluation tasks spanning the domains of lexical, syntactic, and semantic ability, with behavioral data from both children and adults. We evaluate a set of vision–language models on these tasks, comparing models and humans on their response patterns, not their absolute performance. Across tasks, models exhibit variation in their closeness to human response patterns, and models that perform better on a task also more closely resemble human behavioral responses. We also examine the developmental trajectory of OpenCLIP over training, finding that greater training results in closer approximations to adult response patterns. DevBench thus provides a benchmark for comparing models to human language development. These comparisons highlight ways in which model and human language learning processes diverge, providing insight into entry points for improving language models.
@@ 1 Introduction
Recent work has attempted to bridge this data gap by constructing models that are trained on less data, including on naturalistic data from children [2, 3, 4, 5]. However, these models have typically been evaluated using benchmarks that reflect adult human performance, whether explicitly (i.e., with comparison data collected from adult participants) or implicitly (i.e., high accuracy on some recognition task). Such evaluation methods are not appropriate to the goal of understanding whether developmentally realistic data leads to human-like learning. We would not expect child performance to be similar to adult performance on these tasks for various reasons related to both language competence (i.e., children’s vocabulary knowledge) and to their co-developing cognitive abilities (e.g., working memory limitations). Instead, it is crucial to evaluate models on benchmarks that can indicate whether the language ability gained by machine learning models matches the language ability gained by children when exposed to similar developmental data.
@@
Additionally, to conduct methodologically rigorous cognitive evaluation of machine learning models, it is important to compare human and model performance directly (rather than assuming that humans will perform well), and to ensure that evaluation conditions are as similar as possible for models and humans [11, 12, 13]. Notably, many language evaluations with infants and toddlers are multimodal in nature, because this method permits the measurement of language comprehension without the additional working memory and motor control demands of language production. Young children can use nonverbal response methods such as looking or pointing, which can nonetheless reflect their language abilities. The incorporation of multimodality has an additional advantage of introducing grounding, which has been suggested as a possible solution to the data gap [1].
@@
To create a benchmark that satisfies these desiderata, we introduce DevBench, a multimodal benchmark of language evaluation tasks. DevBench includes a suite of seven tasks measuring lexical, syntactic, and semantic ability, with human data from both children and adults. Notably, the primary evaluation metric reflects models’ similarity to human response patterns, rather than absolute performance levels, allowing us to capture finer-grained details about human–model similarity.
@@ 2.1 Multimodal benchmarks
Since the advent of multimodal models, various benchmarks have been developed to evaluate their performance [15]. A majority of these benchmarks primarily involve visual question answering [16, 17, 18, 19, 20, 21]; other tasks include image captioning [18, 22], prediction [23], and retrieval [17, 22].
@@ 2.2 Developmentally inspired evaluation
Some recent work in machine learning evaluation has used benchmarks inspired by developmental psychology and cognitive science. For example, the Large Language Model Response Score (LRS) [24] draws from key experiments in the child development literature to construct question answering tasks for large language models, though it does not involve comparison to data from children. In addition, most of its tasks were originally multimodal (with visual stimuli and verbal prompts), and it is not clear how the translation into a unimodal verbal task would affect human performance. In the visual domain, the Infant-Level Physical Reasoning Benchmark (InfLevel) [25] aimed to evaluate video models on physical reasoning, drawing from classic violation of expectation tasks in the infant cognition literature related to the principles of continuity, solidity, and gravity. However, this benchmark also did not have a direct comparison with human data. The Machine Word Learning (MEWL) benchmark [26] is a multimodal evaluation that builds upon hypothesised mechanisms of few-shot word learning in children, and requires the inference of the meaning of novel words from a set of visual scenes with referentially ambiguous labels. The benchmark contains data from adults, but their poor performance on some subsets of these trials (e.g., those requiring pragmatic implicature) suggests that children might find these tasks very difficult. More recently, the ModelVsBaby benchmark [27] has assessed out of distribution object recognition with corresponding data from 2-year-olds, providing a first step towards direct developmental model–human comparison.
@@ 3 DevBench description
DevBench contains seven tasks across lexical, syntactic, and semantic domains. Each task is accompanied by item-level human data so that full human response distributions can be compared to model scores. The lexical tasks measure vocabulary knowledge, operationalised as the ability to correctly pick out the visual referent of a noun label. The syntactic tasks measure grammatical knowledge, operationalised as the ability to correctly pick out a scene containing the correct relations among its constituent members when given a sentential label. The semantic tasks measure conceptual (visual or linguistic) representation space via representational similarity. A schematic of the tasks and the ages of participants contributing data for each task is shown in Figure 1, and sample trials are shown in Figure 2. All code and data are available at github.com/alvinwmtan/dev-bench; all data are licensed under CC BY-NC-SA or more permissible licences, and have been anonymised to eliminate any personally identifiable information.
@@ 3.1 Tasks
In this task, children are presented with two visual images (one target and one distractor) as well as a verbal cue (e.g., “Look at the dog!”) [32]. Children’s proportion looking time to each image is measured. We used data from three datasets to improve age coverage: 18-month-old data from Adams et al. [33], 24-month-old data from Frank et al. [34], and 30-month-old data from Donnelly and Kidd [35] (total $N$ = 294). Some images in the original stimuli set were not shareable due to license restrictions; in these cases, we used replacement stimuli matched for visual and semantic properties.
@@
In this task, participants were presented with a cue word, and asked to name the first word association that came to mind. We use association frequencies as a proxy for similarity scores. We combined a child dataset from Entwisle [41] with a subset of the adult dataset from Nelson et al. [42] comprising all cue words found in the child dataset (total $N$ > 7040), and thresholded the included responses to omit idiosyncratic responses.
@@
In this task, 4-, 10-, and 14-month-old infants saw pairs of images, and proportion looking times to each image was measured [43] (total $N$ = 73). We used replacement stimuli for some images that were not shareable in the original stimuli set. Dissimilarities between images were calculated as the difference in proportion looking times to the two images.
@@ 3.2 Human–model comparison
Our goal on the lexicon and syntax tasks was to compare the distribution of human choices across response options to model choices. We obtained image–text matching logits for each response option, and calculated the softmax-optimised Kullback–Leibler divergence of human responses from model responses. This novel metric was operationalised as the minimum KL divergence between the human response probability distributions $\mathbf{h}$ from model logits $\mathbf{m}$, optimizing the softmax exponent $\beta$. We average this divergence across trials $t$, and calculate each distribution across images $i$:
@@
For WAT, we calculated the softmax-optimised KL divergence of human association probabilities from softmaxed model text embedding similarities, averaged over all trials. For the visual semantic tasks, we instead conducted human–model comparison by applying representational similarity analysis (RSA) [47] on human and model representational similarity matrices, which represents correlations in the representational geometries of humans and models. All comparisons were conducted within each age bin when applicable.
Our method applies to models from which image–text matching scores could be directly extracted (i.e., similarity models). More recent models have alternatively integrated visual and language inputs via conditional text generation (i.e., generation models). There is as yet limited consensus for the best method to obtain image–text matching scores for generation models; we conducted two exploratory evaluation methods relying on next-token prediction and on log-likelihood measurement. More details on the evaluation of generation models can be found in Appendix C.
3.3 Baselines
We also constructed two baselines for the benchmark to demonstrate the dynamic range that is possible for each task. First, we calculated a human baseline for tasks on which participant-level data were available (LWL, VV, TROG, VOC). To estimate this baseline, we randomly split the participants into two groups and calculated the between-group softmax-optimised KL divergence or RSA similarity as appropriate, repeating for 1000 random splits. We used the median result as a point estimate of the human baseline, which serves as a positive baseline for our tasks. Note that this baseline is an underestimate of the true values, because the results are not corrected upwards for the attenuation due to splitting the data in half.
We also added a random baseline for all tasks, generated using a random initialisation of the OpenCLIP model. The random baseline serves as a negative baseline for our tasks.
@@ 4 Benchmark
We evaluated a diverse set of vision–language models on our benchmark, using an NVIDIA T4 GPU, an NVIDIA A40 GPU, or CPUs (depending on resource availability). A summary of model performance for each task (averaged across all ages) is shown in Table 2, along with model characteristics (number of parameters and size of training set). For lexicon and syntax tasks as well as WAT, we report $D^{*}_{\textup{KL}}$, for which lower scores indicate greater human-likeness in response patterns. For visual semantic tasks (VOC, THINGS), we report RSA similarity, for which higher scores indicate greater human-likeness in response patterns. Divergences are not comparable across datasets due to different features of each dataset. A description of all models and more details on the evaluation setup can be found in Appendix A.
Table 2: Model characteristics and performance across all tasks, demonstrating variation across models. Arrows indicate the direction of better performance (i.e., lower is better vs. higher is better). Bolded results indicate most human-like result on a task.
@@
Human 0.010 0.091 0.028 0.251
Random (OpenCLIP) 1.0B 0 0.087 0.740 0.908 0.258 0.495 0.246 0.054
Overall, CLIP-large and OpenCLIP-H performed relatively well on the lexicon and syntax tasks. CVCL, which was trained on a small set of annotated head-mounted camera from infants [4], was mostly dissimilar to humans in lexicon and syntax tasks, even on the looking-while-listening (LWL) task, which was administered to young children. However, CVCL outperformed other models on the visual object categorization task, suggesting that its visual representational space was more similar to infants’. SigLIP was also unusually poor-performing (especially given its general performance and accuracy), potentially suggesting that aspects of its training (e.g., its objective function) may have resulted in non-alignment with humans. We thus excluded it as an outlier from further analyses.
@@ 6.1 Limitations and future work
The development of DevBench was constrained by currently available human data; notably, this means that some tasks have relatively few items (in comparison with many other machine learning benchmarks), and some tasks have relatively few human participants. These limitations may result in uncertain reliabilities for the obtained results. More data are currently being collected for some of the tasks in DevBench, which we anticipate will improve reliability, and simultaneously permit more expansive developmental analyses due to an extension of the included age ranges. However, we hope that this approach and standardized format will enable future researchers to contribute novel datasets to DevBench, which we anticipate may grow in the coming years. Indeed, DevBench also includes primarily tasks for English-speaking children and adults, due to dataset availability. This situation reflects inequalities in language acquisition research [56], and more data is needed to construct a multilingual version of DevBench, which will be more comprehensive and generalisable.
Additionally, model performance in our evaluation setup may be affected by the domain gap between models’ training data and the stimuli used in our benchmark; for example, TROG uses cartoon depictions of events, which are dissimilar to the more photorealistic training data of CLIP. Thus, our evaluation results likely represent a lower bound on model–human similarity. Children as young as two years of age are able to learn from and generalise to pictographic depictions of objects [57, 58, 59], however, suggesting that generalisation across representations is an early-acquired skill.
In our analyses of the relationship between model–human similarity and model features, we could not hold model architectures constant due to the limited availability of relevant model checkpoints – thus we cannot make strong claims about precisely what aspects of models lead to better fit to human responses. Systematic evaluation of the roles of training data and model size thus remains an important research question for future work (see [60, 61] for related work on scaling).
@@ 6.2 Conclusion
We hope that DevBench can serve as an encouragement for machine learning researchers and cognitive scientists to develop vision–language models that not only perform well, but also can more closely approximate human learning. In particular, DevBench highlights the need for more fully open models with training checkpoints [63], enabling the study of training trajectories, as well as the need for more human-realistic training [2, 60] to better characterise model–human correspondences across developmental change. It may also be possible to adopt a developmental, multimodal approach to studying domains other than language, such as mathematical, logical, and social reasoning; more intentional data collection across a wide range of ages, tasks, and contexts will help to provide an increasingly comprehensive set of comparison data to better understand model learning trajectories (e.g., [64]). These research directions, among others, will help us to better understand the processes underlying human development, and how we might transfer humans’ learning efficiencies onto machine learning models.
@@
Appendix A Additional method details
A.1 Additional task details
A breakdown of the demographics for each age bin for each task is shown in Table 4. An overview of key methodological dimensions for the sources of each task is shown in Table 5.
Table 4: Participant demographics for all tasks. A: Adult.
Task Age bin Mean age N Country of origin Test language
LWL 1.5 1.42 166 USA English
2 2.49 28 USA English
2.5 2.56 100 Australia English
VV 4 5.5 178 USA English
7 8.22 458 USA English
10 10.17 939 USA English
A 205 USA English
TROG 11 11.77 514 USA English
WG A 171 Not specified English
WAT 5 5.6 200 USA English
6 6.55 280 USA English
8 8.73 280 USA English
10 10.45 280 USA English
A 6000 USA English
VOC 0.3 0.33 24 France French
0.8 0.9 24 France French
1.6 1.6 25 France French
THINGS A 12340 Not specified English
Table 5: Overview of selected methodological dimensions for all tasks. A: Adult.
Task Source Age bins Recruitment
method
Administration
method
Data collection
method
Experimental
setup
Primary goal of study
LWL Adams et al. 1.5 Hospital
and others
In-person Eyetracking Preferential
looking
Investigate associations between caregiver talk and language skills in full/preterm children
Frank et al. 2 Museum In-person Eyetracking Preferential
looking
Investigate the reliability of using tablets to collect experimental data from young children
Donnelly & Kidd 2.5 Not specified In-person Eyetracking Preferential
looking
Investigate the relationship between phonological onset density and lexical access
VV Long et al. 4, 7, 11, A Schools
and online
In-person
and online
4AFC Best match Investigate developmental changes in visual concept knowledge
TROG Silverman & Yeatman 11 Schools In-person 4AFC Best match Investigate developmental changes in grammatical knowledge
WG Thrush et al. A Amazon
MTurk
Online 2AFC Match or
non-match
Investigate visuo–linguistic compositionality in vision–language models
WAT Entwisle 5, 6, 8, 10 Schools In-person Oral
response
Word
association
Investigate chidlren’s verbal conceptual space and its relationship with demographic factors
Nelson et al. A Not specified In-person Written
response
Word
association
Investigate adults’ verbal conceptual space and cue-to-target association strengths
VOC Spriet et al. 0.3, 0.8, 1.6 Not specified In-person Eyetracking Preferential
looking
Investigate infants’ visual conceptual space and its relationship with adult brain activity
THINGS Hebart et al. A Amazon
MTurk
Online 3AFC Odd one out Investigate adults’ visual conceptual space and its relationship with adult brain activity
A.2 Evaluated models
@@
SigLIP
SigLIP [54] is a modification of CLIP which does not rely on batch-wise normalisation, using a sigmoid loss instead of a softmax loss. It uses a ViT vision encoder and a transformer text encoder, and treats each image–text pair as a binary classification problem (match vs non-match).
@@
A.3 Evaluation setup
For lexicon and syntax tasks, models were evaluated by passing in each image–text pair for each trial as inputs, and obtaining model logits for each pair. For LWL and WG, there were two images in each trial, while for VV and TROG, there were four images in each trial. Model logits were then used to calculate the softmax-optimised KL divergence with human responses.
For VOC and THINGS, we obtained image embeddings for each stimulus, and obtained a representational similarity matrix (RSM) by calculating the pairwise cosine similarity for each pair of images. We then compared the model RSM with that obtained from human responses by calculating the Spearman’s rank correlation coefficient for entries below the main diagonal in model and human RSMs.
For WAT, we obtained text embeddings for each stimulus, and calculated the pairwise cosine similarity for all cue–target pairs in the human response data. Model similarity values were then used to calculate the softmax-optimised KL divergence with human response distributions for each cue word.
Some models (e.g., BridgeTower) always required both image and text inputs. For these models, we used an empty string as the dummy text input when obtaining image embeddings for VOC and THINGS, and we used a neutral gray square as the dummy image input when obtaining text embeddings for WAT.
A.4 Softmax-optimised Kullback–Leibler divergence
We used the softmax-optimised KL divergence as our novel metric of model–human dissimilarity. (Ordinary) KL divergence reflects how different a target probability distribution is from a reference probability distribution, often considered the true distribution. In the case of DevBench, the reference distribution is obtained from human responses, while the target distribution is obtained from model responses.
The typical method of deriving probabilities from model responses is by conducting a softmax over logits. However, we considered that model logits may not be calibrated to the same scale as human responses, and therefore included the temperature, $\beta$, as a free parameter. In other words, the resultant distribution after optimisation can be considered a one-parameter projection from logit space to probability space, and the best fitting projection is that which induces the minimum KL divergence to the human response distribution.
Appendix B Additional detailed results
B.1 Detailed benchmark results